What is the area enclosed by a circular sector whose radius is r and arc length is s. : What is the area enclosed by a circular sector whose radius is r and arc length is s.
Answer by Earlsdon(3748)

You could try proportions on this problem:
Let's compare the ratio of the arc length, S, to the circumference of the circle, C, with the ratio of the area enclosed by the sector A[s] to the area of the entire circle A[c] = (pi)r^2. Remember that C = 2(pi)r
S/C = A[s]/A[c]
S/2(pi)r = A[s]/(pi)r^2 Simplify and solve for A[s]
S(pi)r^2/2(pi)r = A[s]
A[s] = Sr/2
Question 149138: What is the area enclosed by a circular sector whose radius is r and arc length is s. : What is the area enclosed by a circular sector whose radius is r and arc length is s.
Answer by edjones(2401)

pi*r^2=A (circle)
pi*2r=C (circle)
C/s=2pi radians/s [s is in radians]
A {circle}*(s/(2pi radians))=A {enclosed by the circular sector}
.
Ed